# The Probabilistic Foundations of Surveillance Failure: From False Alerts to Structural Bias

### v1

@@ 2.4 Sharp Threshold Behavior
The transition occurs in a window $\Delta\alpha\sim 1/(\lambda D(c\|1))\to 0$, becoming arbitrarily sharp as $\lambda$ increases. This follows from standard large-deviation theory for sums of i.i.d. rare events (Dembo and Zeitouni, 2010).
Proof.
From Theorem 2.4, we have the two-sided bounds
which together imply $\log q=-\lambda D(c\|1)+O(\log\lambda)$ as $\lambda\to\infty$ (since $q\geq C\lambda^{-1/2}e^{-\lambda D}$ and $q\leq e^{-\lambda D}$, taking logarithms gives the stated error term). With $n=\sqrt{\lambda}\,e^{\alpha\lambda D(c\|1)}$, we have
$\log(nq)=\log n+\log q=\frac{1}{2}\log\lambda+\alpha\lambda D(c\|1)-\lambda D(c\|1)+O(\log\lambda),$
which simplifies to
@@
Hence for $\alpha<1$, $nq\to 0$ exponentially; for $\alpha>1$, $nq\to\infty$ exponentially. The transition width $\Delta\alpha$ satisfying $|(\alpha-1)\lambda D|=O(1)$ gives $\Delta\alpha=O(1/(\lambda D))$, which vanishes as $\lambda\to\infty$. Equivalently, the absolute window in $\log n$ is $O(1)$, making the transition arbitrarily sharp in the exponential scale. The limiting behavior (5) follows from $\Pr(\text{false alert})\approx 1-e^{-nq}$, which approaches 0 when $nq\to 0$ and approaches 1 when $nq\to\infty$. ∎
@@
This sharp threshold implies no gradual degradation: systems operating near the critical line $\log n=\lambda D(c\|1)$ transition from reliable to unreliable essentially instantaneously as $\lambda$ increases (Figure 1). Combined with exponential data growth $\lambda(t)=\lambda_{0}\gamma^{t}$, every system inevitably crosses this threshold at finite time $T^{*}\sim\log(m/\lambda_{0})/\log\gamma$ (Theorem 3.2), making failure predictable rather than merely possible.
3 Temporal Dynamics: Finite System Lifetimes
@@
In the preceding analysis (Section 2), we studied thresholds of the form $m=c\lambda$ with $c>1$ constant, where both $m$ and $\lambda$ scale together. In this section, we adopt a different perspective: we fix the threshold $m$ at deployment time while allowing $\lambda(t)=k(t)\cdot p$ to grow over time as data accumulates. This models realistic system operation, where detection thresholds are predetermined policy parameters that remain constant even as surveillance capabilities expand. The critical behavior occurs when the growing $\lambda(t)$ crosses the fixed threshold $m$, at which point false alerts become inevitable.
3.1 Exponential Data Growth
Modern data collection exhibits exponential growth (Hilbert and López, 2011). We model this as:
@@ 3.1 Exponential Data Growth
3.2 Critical Time Derivation
@@
For populations large enough that system failure occurs near $\lambda(T^{*})\approx m$ (rather than deep in the Poisson tail), the system becomes unreliable at time
@@ 3.3 Key Insights
    Finite lifetime is inevitable: For any fixed threshold $m$ and exponential growth $\gamma>1$, we have $T^{*}<\infty$. Even if one attempted to maintain reliability by scaling the threshold with data ($m\propto\lambda(t)$), this would require the monitored population to grow as $n_{\text{crit}}(t)\sim\exp(c\gamma^{t})$, doubly-exponential growth that far exceeds realistic population dynamics. Real surveillance systems use fixed thresholds and monitor approximately fixed populations, making temporal failure inevitable. [This mathematical structure parallels the “double birthday paradox” analysis of Pollanen (2024), though arising here in a distinct screening context.]
@@ 3.5 Practical Implications
Figure 1: Phase transitions in surveillance system reliability. Systems exhibit sharp transitions from reliable to unreliable operation across four dimensions: (a) attribute growth with Monte Carlo validation (orange points, 5000 runs) confirming theoretical predictions (mean error << 0.002), (b) population scaling (log $n$ vs. false-alert probability), (c) temporal dynamics ($t$ vs. false-alert probability under exponential data growth), and (d) group dominance showing differential exposure creates exponential false-alert disparities. All transitions sharpen as system parameters increase, consistent with large-deviation theory.
4 Heterogeneous Population Structure
Not all individuals have equal representation in surveillance databases. Socioeconomic, geographic, and demographic factors lead to differential exposure rates (Lyon, 2003; Eubanks, 2018).
4.1 Multi-Group Model
@@ 4.2 Dominance and Disparity
Let $g^{*}=\arg\max_{g}\{n_{g}q_{g}\}$ be the group contributing the most to system risk. Then:
@@
4.3 Fairness Implications
The Group Dominance Effect (Theorem 4.2) has important implications for fairness in surveillance systems. When different population groups experience differential surveillance exposure, small differences in exposure rates create exponential disparities in outcomes. Proposition 4.3 shows that exposure ratios of 2–4 times generate false alert disparities exceeding 20 times near critical thresholds, due to Poisson tail behavior. This exponential amplification means that even modest differences in surveillance intensity produce severe outcome inequalities.
Proposition 4.3 (Exposure Amplification Through Poisson Tails).
  1. 1.
    $\frac{n_{2}q_{2}}{n_{1}q_{1}}=\frac{q_{2}}{q_{1}}\geq\exp(c\cdot m)$ (11)
    for some constant $c>0$ (depending on $\lambda_{1}$), for all sufficiently large $m$. (Since $q_{1}\lesssim(\lambda_{1}/m)^{m}$ for $m>\lambda_{1}$, while $q_{2}\geq 1/2-O(m^{-1/2})$ by the central limit theorem, the ratio $q_{2}/q_{1}$ grows super-exponentially in $m$. In particular, for any fixed constant $c>0$ there exists $M(c)$ such that for all $m\geq M(c)$,
    $\frac{q_{2}}{q_{1}}\;\geq\;\exp(cm).$
    Thus the disparity eventually dominates $\exp(cm)$ for arbitrary but fixed $c>0$.) Indeed, $q_{2}\geq\Pr(\text{Poisson}(m)\geq m)$ by monotonicity, and for large $m$, $\Pr(\text{Poisson}(m)\geq m)\geq 1/2-O(m^{-1/2})$, while $q_{1}\leq\exp(-\lambda_{1}D(m/\lambda_{1}\|1))$ by the Chernoff bound in Lemma 2.3, which is exponentially small for $m>\lambda_{1}$.
  2. 2.
  3. 3.
Proof.
(1) When $\lambda_{2}\geq m$, by monotonicity of the Poisson tail probability in the parameter $\lambda$, we have
For large $m$, by symmetry of the limiting normal approximation, $\Pr(\text{Poisson}(m)\geq m)$ approaches $1/2$ from below. The median of the Poisson($m$) distribution satisfies $m-\log 2\leq\text{median}\leq m+1/3$, so the tail probability at $m$ is bounded below by $1/2-O(m^{-1/2})$. Meanwhile, $q_{1}\leq\exp(-\lambda_{1}D(m/\lambda_{1}\|1))$ by Lemma 2.3, which is exponentially small when $m>\lambda_{1}$. The ratio $q_{2}/q_{1}$ therefore grows at least exponentially in $m$: for some constant $c>0$ (depending on $\lambda_{1}$), we have $q_{2}/q_{1}\geq\exp(c\cdot m)$ for sufficiently large $m$.
(2) Follows from (1) by noting $q_{2}/(q_{1}+q_{2})\to 1$ when $q_{2}/q_{1}\to\infty$.
(3) Group 2 reaches $\lambda_{2}=m$ when $k=m/p_{2}$, while Group 1 reaches $\lambda_{1}=m$ when $k=m/p_{1}>m/p_{2}$. The disparity is maximal in the interval $m/p_{2}<k<m/p_{1}$. ∎
Remark 4.4 (Structural vs. Algorithmic Bias).
The exponential amplification in part (1) is particularly striking: when both groups are screened using the same attribute set and threshold, small differences in exposure translate to exponential differences in false alert rates (Figure 1(d)). The effect manifests temporally as different groups reach critical false-alert rates at different times: Group 2 with exposure rate $p_{2}=0.020$ fails at $k_{2}^{*}=350$ attributes, while Group 1 with $p_{1}=0.005$ remains reliable until $k_{1}^{*}=1400$. This fourfold difference in system lifetime arises from the fourfold difference in exposure rates, but the amplification becomes exponential when comparing simultaneous false alert rates: at intermediate $k\in(350,1400)$, Group 2 experiences exponentially more false alerts than Group 1.
Our exponential disparity result formalizes concerns from the algorithmic fairness literature about the inadequacy of threshold-based fairness metrics (Hardt et al., 2016; Barocas et al., 2019). Even attempting to equalize false positive rates through group-specific threshold adjustments cannot eliminate outcome disparities, as the required thresholds themselves encode the underlying exposure inequalities. This suggests fairness cannot be achieved through algorithmic tuning alone (Dwork et al., 2012), but requires equalizing data collection intensity at the source. This conclusion extends recent work on “fairness through awareness” and formalizes “structural bias” arguments (Eubanks, 2018; Benjamin, 2019) showing that disparities are intrinsic to systems built on heterogeneous data collection, not merely implementation artifacts.
4.4 Numerical Example: Geographic Disparities

### v2

@@ 3.1.4 Sharp Threshold Behavior
The transition occurs in a window $\Delta\alpha\asymp 1/(\lambda D(c\|1))\to 0$, becoming arbitrarily sharp as $\lambda$ increases; this follows from standard large-deviation theory for sums of i.i.d. rare events [29]. (Note: $\alpha$ here denotes the population scaling exponent, distinct from the posterior probability threshold $\alpha$ in Section 4.1.)
Proof sketch.
From Theorem 3.5, $\log q=-\lambda D(c\|1)+O(\log\lambda)$. With $n=\sqrt{\lambda}\,e^{\alpha\lambda D(c\|1)}$, we obtain
@@
For $\alpha<1$, $nq\to 0$ exponentially; for $\alpha>1$, $nq\to\infty$ exponentially. The transition width $\Delta\alpha=O(1/(\lambda D))\to 0$. Full details appear in Appendix A.2. ∎
@@
This sharp threshold implies no gradual degradation: systems operating near the critical line $\log n=\lambda D(c\|1)$ transition from reliable to unreliable essentially instantaneously as $\lambda$ increases (Figure 1). Combined with exponential data growth $\lambda(t)=\lambda_{0}\gamma^{t}$ with $\lambda_{0}=k_{0}p$, every system inevitably crosses this threshold at a finite time
$T^{*}\sim\frac{\log(m/\lambda_{0})}{\log\gamma}=\frac{1}{\log\gamma}\log\!\left(\frac{m}{k_{0}p}\right)$
(Theorem 3.9), making failure predictable rather than merely possible.
3.1.5 Numerical Validation
Monte Carlo simulations (5000 runs per data point) validate the Poisson approximation and critical population predictions as $k$ increases. Figure 1(a) compares simulated false-alert rates (orange points) to theoretical predictions, showing mean absolute error of 0.0015 across the critical transition region. The sharp threshold behavior predicted by Proposition 3.6 is clearly evident in the simulation data.
The accuracy of the Poisson approximation across varying $(k,p)$ combinations follows from Le Cam’s inequality (Section 3.1.1). In the homogeneous case $p_{i}=p$ used in this section, Le Cam’s bound simplifies to
$d_{\mathrm{TV}}\leq 2kp^{2}=2\lambda p,$
[…]
Table 1: Poisson approximation error: theoretical bound versus actual total variation distance for representative $(k,p)$ combinations.
$k$ $p$ $\lambda=kp$ Le Cam Bound Actual $d_{\mathrm{TV}}$
50 0.10 5.0 1.000 0.026
100 0.05 5.0 0.500 0.013
500 0.01 5.0 0.100 0.0025
1000 0.005 5.0 0.050 0.0012
5000 0.001 5.0 0.010 0.0002
100 0.10 10.0 2.000 0.026
1000 0.01 10.0 0.200 0.0025
The Le Cam bound is conservative by a factor of 17–40 in these regimes; actual approximation error is substantially smaller than the bound suggests. For the primary parameters used throughout this paper ($k=1000$, $p=0.005$, $\lambda=5$), the actual total variation distance is approximately $0.001$, indicating that the Poisson approximation introduces negligible error relative to the phase transition effects we characterize. Even in less favorable regimes ($k=50$, $p=0.1$), the approximation error remains below $0.03$, confirming that the Poisson model is appropriate across the full range of parameters considered.
Additional numerical illustrations across a broader range of parameters, as well as examples based on real-world datasets, are given in Appendix C.
Key Takeaway. The per-person false alert probability decays as $q\sim e^{-\lambda D(c\|1)}$, and the system becomes unreliable once the population size exceeds $n_{\mathrm{crit}}\sim\sqrt{\lambda}\,e^{\lambda D(c\|1)}$. This behavior produces a sharp phase transition: systems move abruptly from reliable to unreliable operation, and the sharpness of the transition increases monotonically with $\lambda$.
3.2 Temporal Dynamics: Finite System Lifetimes
@@
In the preceding analysis (Section 3.1), we studied thresholds of the form $m=c\lambda$ with $c>1$ constant, where both $m$ and $\lambda$ scale together. In this section, we adopt a different perspective: we fix the threshold $m$ at deployment time while allowing $\lambda(t)=k(t)\cdot p$ to grow over time as data accumulates. This models realistic system operation, where detection thresholds are predetermined policy parameters that remain constant even as surveillance capabilities expand. The critical behavior occurs when the growing $\lambda(t)$ crosses the fixed threshold $m$, at which point false alerts become inevitable.
3.2.1 Exponential Data Growth
Modern data collection exhibits exponential growth [30, 31]. We model this as:
@@ 3.2.1 Exponential Data Growth
3.2.2 Critical Time Derivation
@@
For populations satisfying $n\gtrsim\sqrt{m}\exp(m/2)$ (such that system failure occurs near $\lambda(T^{*})\approx m$ rather than deep in the Poisson tail where $\lambda\ll m$), the system becomes unreliable at time
@@ 3.2.3 Key Insights
    Finite lifetime is inevitable: For any fixed threshold $m$ and exponential growth $\gamma>1$, we have $T^{*}<\infty$.
    Even if one attempts to preserve reliability by scaling the threshold with the data ($m\propto\lambda(t)$), doing so would require the monitored population to grow as $n_{\mathrm{crit}}(t)\sim\exp(C\gamma^{t})$ for some C>0C>0—a doubly-exponential rate that far exceeds any realistic population dynamics.
    Because real surveillance systems use fixed thresholds and monitor approximately fixed populations, temporal failure becomes inevitable rather than merely possible. This mathematical structure parallels the “double birthday paradox” in our earlier work [32], though it arises here in a distinct screening context.
@@ 3.2.5 Practical Implications
Remark 3.11 (Non-Exponential Data Growth).
The analysis above focuses on exponential growth, but alternative data-growth models lead to qualitatively different system lifetimes. For simplicity, we identify the onset of system unreliability with the heuristic condition $\lambda(t)=p\,k(t)\approx m$, which corresponds to the point at which false-alert probabilities increase sharply. We assume $m>pk_{0}$ so that this regime is reached, if at all, at some $T^{*}>0$.
  • Polynomial growth $k(t)=k_{0}+\gamma t^{\alpha}$ with $\alpha\geq 1$: Solving $p(k_{0}+\gamma T^{*\alpha})\approx m$ gives
    $T^{*}\approx\left(\frac{m/p-k_{0}}{\gamma}\right)^{1/\alpha},$
    so $T^{*}\propto m^{1/\alpha}$ grows linearly when $\alpha=1$ and sublinearly when $\alpha>1$. Larger values of $\alpha$ correspond to faster growth of $k(t)$, leading to shorter system lifetimes. Moreover, because $1/\alpha$ decreases with $\alpha$, increases in $m$ yield progressively smaller gains in $T^{*}$, further reducing threshold protection. Nonetheless, polynomial growth still provides significantly more protection than exponential growth, where $T^{*}$ increases only logarithmically with $m$.
  • Logarithmic growth $k(t)=k_{0}+\beta\log(1+\gamma t)$: Here
    $T^{*}\approx\frac{\exp\!\big((m/p-k_{0})/\beta\big)-1}{\gamma},$
    so the time required for $\lambda(t)$ to reach $m$ grows exponentially in the threshold $m$. Because $k(t)$ itself grows only logarithmically in $t$, this time can be extremely large, and practical data collection may saturate before a critical regime is ever reached.
  • Bounded growth $k(t)\to k_{\max}$: If the maximal attainable rate $\lambda_{\max}=p\,k_{\max}$ satisfies $\lambda_{\max}<m$, the system never enters the high-false-alert regime and remains below the critical level for all time.
Exponential growth is the critical case because it eventually exceeds any subexponential threshold-adjustment strategy: if $\lambda(t)$ grows exponentially while $m(t)$ increases subexponentially, then $\lambda(t)>m(t)$ for all sufficiently large $t$. Empirical studies suggest that historical data-collection capacity has grown approximately exponentially [30], making this the scenario of greatest relevance for policy analysis.
Figure 1: Phase transitions in surveillance system reliability. Systems exhibit sharp transitions from reliable to unreliable operation across four dimensions. (a) Attribute growth: Monte Carlo simulations (orange points, 5000 runs per data point) validate the theoretical predictions (blue curve) with mean absolute error below 0.002. Takeaway: The sharp S-curve illustrates a phase transition: systems remain reliable until a critical attribute count is reached, after which reliability collapses rapidly with almost no intermediate zone. (b) Population scaling: False-alert probability as a function of population size (log scale) for fixed $\lambda$ and threshold. Takeaway: The transition sharpens as $\lambda$ increases, confirming Proposition 3.6. Populations below $n_{\mathrm{crit}}\sim e^{\lambda D}$ are reliable, while those above this scale almost certainly generate false alerts. (c) Temporal dynamics: Under exponential data growth with $\gamma=1.5$ (50% annual growth), systems transition from reliable to unreliable at a predictable time $T^{*}\approx 4$ years (Theorem 3.9). Takeaway: Degradation is abrupt rather than gradual—systems remain functional until they cross a critical time threshold and then fail rapidly. (d) Group dominance: Two groups with different exposure rates ($p_{1}=0.005$ vs. $p_{2}=0.02$) exhibit markedly different false-alert trajectories. Takeaway: The high-exposure group (solid) reaches the failure regime much sooner than the low-exposure group (dashed), illustrating the structural exposure-driven disparities described in Proposition 3.15.
Key Takeaway. Under exponential data growth, every screening system has a finite, calculable lifetime $T^{*}\approx\frac{1}{\log\gamma}\log(m/k_{0}p)$ determined by when $\lambda(t)$ reaches the threshold $m$. Threshold adjustments provide only logarithmic protection against exponential growth; the growth rate $\gamma$ dominates system longevity, making temporal failure inevitable rather than merely possible.
3.3 Heterogeneous Population Structure
Not all individuals have equal representation in surveillance databases. Socioeconomic, geographic, and demographic factors lead to differential exposure rates [5, 33].
3.3.1 Multi-Group Model
@@ 3.3.2 Dominance and Disparity
Let $g^{*}=\operatorname*{arg\,max}_{g}\{n_{g}q_{g}\}$ be the group contributing the most to system risk. Then:
@@
3.3.3 Numerical Example
Example 3.14 (Geographic Disparities).
@@ 3.3.3 Numerical Example
In the UCI Adult Census data, increasing the threshold produces the predicted divergence in group contributions: at $m=8$, a 2% subgroup generates 17% of false alerts, while the lowest-exposure group contributes none. See Appendix C.
3.3.4 Exposure Amplification Through Poisson Tails
Proposition 3.15 (Exposure Amplification Through Poisson Tails).
  1. 1.
    $\frac{n_{2}q_{2}}{n_{1}q_{1}}=\frac{q_{2}}{q_{1}}\geq\exp(c\cdot m)$ (11)
    for some constant $c>0$ depending on $\lambda_{1}$, for all sufficiently large $m$. The disparity grows super-exponentially because $q_{1}$ decays exponentially while $q_{2}$ remains bounded away from zero.
  2. 2.
  3. 3.
Proof sketch.
(1) For $\lambda_{2}\geq m$, we have $q_{2}\geq\Pr(\mathrm{Poisson}(m)\geq m)\geq 1/2-O(m^{-1/2})$ by normal approximation, while $q_{1}\leq e^{-\lambda_{1}D(m/\lambda_{1}\|1)}$ is exponentially small (Lemma 3.3). The ratio grows super-exponentially. (2) Since $q_{2}/q_{1}\to\infty$, we have $q_{2}/(q_{1}+q_{2})\to 1$. (3) Group 2 reaches $\lambda_{2}=m$ at $k=m/p_{2}<m/p_{1}$, so disparity is maximal for $k\in(m/p_{2},m/p_{1})$. Full details appear in Appendix A.3. ∎
Remark 3.16 (Within-Group Individual Heterogeneity).
Even within a demographic group $g$, individuals may experience different exposure levels due to occupation, neighborhood, behavioral patterns, or daily routines, leading to person-specific match probabilities $p_{g}^{(i)}$. A convenient model treats $p_{g}^{(i)}$ as random with mean $\bar{p}_{g}$, for example $p_{g}^{(i)}\sim\mathrm{Beta}(\alpha_{g},\beta_{g})$ with $\mathbb{E}[p_{g}^{(i)}]=\bar{p}_{g}$. This induces a mixture of Poisson distributions for match counts whose variance exceeds the Poisson variance of the homogeneous case (classical overdispersion).
The qualitative phenomena we describe persist under within-group heterogeneity. As long as the group-level means $\bar{p}_{g}$ differ, the higher-exposure group reaches false-alert thresholds earlier on average. Moreover, within-group heterogeneity may actually increase disparity: overdispersion inflates upper-tail probabilities, making heterogeneous groups more likely to produce false alerts than homogeneous groups with the same mean exposure. A full mixture analysis is beyond our present scope, but the structural exposure-disparity mechanism is robust to this generalization.
Key Takeaway. Differential surveillance exposure creates exponential disparities in false-alert rates through Poisson tail behavior. Because this mechanism arises from data collection intensity rather than classifier design, neither threshold adjustment nor standard algorithmic fairness interventions can eliminate it. The disparity persists under within-group heterogeneity and may even intensify due to overdispersion. Achieving genuine outcome parity requires equalizing surveillance intensity itself.
3.4 Effective Dimensionality Under Correlation
Real surveillance data exhibit spatial and temporal dependencies that fundamentally alter tail probabilities. While correlation does not change the expected number of matches ($\mathbb{E}[\sum_{i}X_{i}]=kp$ regardless of dependence), it induces overdispersion that inflates tail probabilities beyond the independent Poisson approximation.
The key insight is that positive correlation reduces the effective degrees of freedom. Standard design-effect methodology parameterizes this through an effective sample size $k_{\mathrm{eff}}$ satisfying $\mathrm{Var}(Y)=k_{\mathrm{eff}}\cdot p(1-p)$. For spatial correlation with exponential decay over area $A$ with correlation length $\xi$:
$k_{\mathrm{eff}}\approx\frac{A}{2\pi\xi^{2}}.$ (12)
For temporal correlation with correlation time $\tau$:
$k_{\mathrm{eff}}\approx\frac{k}{2\tau}.$ (13)
This modifies the critical population scale. The exponent $\lambda D(c\|1)$ is reduced by the factor $k_{\mathrm{eff}}/k$, yielding:
$n_{\mathrm{crit}}^{\mathrm{corr}}\lesssim\sqrt{\lambda}\exp\!\left(\frac{k_{\mathrm{eff}}}{k}\cdot\lambda D(c\|1)\right).$ (14)
The practical impact is severe. City-scale location monitoring with k=10,000k=10{,}000 cells and correlation length $\xi=500$m yields $k_{\mathrm{eff}}\approx 64$, reducing critical populations by over two orders of magnitude. Year-long daily observations ($k=365$) with $\tau=30$ days yields $k_{\mathrm{eff}}\approx 6$, making reliable surveillance of even small groups challenging.
Remark 3.17 (Scope: Heuristic Approximations).
This analysis uses heuristic approximations rather than rigorous large-deviation theory. The effective degrees of freedom approach captures the dominant effect (variance inflation) but does not constitute a formal theorem. The qualitative conclusion that positive correlation accelerates false-alert saturation is robust and well established in the literature on design effects, effective sample size, and correlation-adjusted multiple testing. Full technical details, derivations, and worked examples appear in Appendix B.
4 Discussion
4.1 Bayesian Posterior Reliability and the Base-Rate Trap
4.1.1 The Bayesian Framework and Positive Predictive Value
We introduce standard notation from diagnostic testing and forensic statistics [2]:
$\pi$ $\equiv\Pr(\text{target})\quad\text{(base rate)},$
By Bayes’ rule, the positive predictive value (PPV) is
$\Pr(\text{target}\mid\text{flag})=\frac{s\,\pi}{s\,\pi+q\,(1-\pi)}.$ (15)
$\mathbb{E}[\#\text{ flags}]\approx r\,s+n\,q.$
Thus, the expected false discovery rate (FDR) is
matching (15) exactly.
Remark 4.1 (Sensitivity Dependence on Threshold).
The analysis above treats sensitivity $s$ as constant, but in practice $s=\Pr(\text{flag}\mid\text{target})$ typically decreases as the threshold $m$ increases: stricter criteria miss a larger fraction of true targets. A simple parametric model capturing this tradeoff is
$s(m)=s_{\max}\exp\!\bigl(-\beta(m-\lambda_{\mathrm{signal}})_{+}\bigr),$ (17)
where $s_{\max}$ is sensitivity at low thresholds, $\lambda_{\mathrm{signal}}>\lambda=kp$ is the expected number of matching attributes for a true target, $\beta>0$ controls the decay rate, and $(x)_{+}=\max(x,0)$. The exponential form reflects the common observation that match-score distributions exhibit approximately exponential tails, though any monotone decreasing form would yield the same qualitative conclusions.
Substituting (17) into the PPV formula (16) reveals competing effects as $m$ increases:
  • False positives decrease: for innocent individuals, $q(m)=\Pr(Y\geq m)$ with $Y\sim\mathrm{Poisson}(\lambda)$ and $\lambda=kp$, and Lemma 3.3 shows that $q(m)$ decays very rapidly once $m>\lambda$ (indeed, faster than any fixed-rate exponential in $m$).
  • True positives decrease: $s(m)$ remains near $s_{\max}$ for $m\leq\lambda_{\mathrm{signal}}$ but then decays exponentially for $m>\lambda_{\mathrm{signal}}$.
These effects create an intermediate regime in which PPV may plateau or improve more slowly than the constant-$s$ analysis suggests, particularly once $m$ exceeds $\lambda_{\mathrm{signal}}$ and losses in sensitivity offset some of the gains from reduced false positives. Under the model (17), however, Lemma 3.3 implies that $q(m)$ eventually decays much faster than $s(m)$, so $q(m)/s(m)\to 0$ and therefore $\mathrm{PPV}(m)\to 1$ as $m\to\infty$. Any plateau or dip in PPV can therefore occur only over this intermediate range of thresholds, not asymptotically.
Because sensitivity is bounded above by $s_{\max}$, the constant-$s$ model used earlier provides an upper bound on achievable PPV for any given threshold: $\mathrm{PPV}(m)$ is increasing in $s$, and $s(m)\leq s_{\max}$ for all $m$. Optimal threshold selection ultimately requires specifying both the innocent and target match distributions—equivalently, full ROC curve analysis [23]. Since the signal distribution is application-dependent, we do not pursue this direction here. Of course, driving $m$ to extremely large values also drives the overall flag rate toward zero, so the $\mathrm{PPV}(m)\to 1$ limit is primarily of theoretical interest; in practice, the operationally relevant regime is the intermediate range of thresholds where nontrivial detection rates are maintained.
4.1.2 Posterior Reliability and Bayesian Critical Scales
Remark 4.2 (Sparse-Target Regime).
The following proposition applies to the “needle in a haystack” setting where $r$ (the expected number of true targets) is fixed or grows slowly while population $n$ increases. If instead $r\propto n$ (constant prevalence), then the ratio $r/n$ is fixed, and the condition for actionable PPV reduces to a bound on $q/s$ relative to $\alpha/(1-\alpha)$, independent of $n$. The sparse-target regime is the most challenging case for screening systems and is therefore the primary focus of this analysis.
Proposition 4.3 (Bayesian Critical Population for Actionable PPV).
$\Pr(\text{target}\mid\text{flag})\geq\alpha$
is satisfied whenever
$n\leq\frac{(1-\alpha)\,r\,s}{\alpha\,q}.$ (18)
Proof.
When sensitivity varies with threshold as in Remark 4.1, the bound (19) should be interpreted as an upper bound on achievable Bayesian reliability, since $s(m)\leq s_{\max}$ for all $m$.
Remark 4.4 (The Bayesian Trap).
4.1.3 Likelihood Ratios and Classical Fallacies
$L=\frac{s}{q},$
and mistakenly infer that $\Pr(\text{target}\mid\text{flag})$ is therefore large. This is the classical prosecutor’s fallacy [2]. Bayes’ rule shows:
4.1.4 Resolving the DNA Database Controversy
Forensic statisticians have long debated the evidential value of DNA “cold hits.” Stockmarr [1] argued that searching databases weakens evidence by inflating coincidental match probabilities; Balding [34, 2] countered that likelihood ratios preserve evidential weight. Our framework resolves this apparent contradiction by identifying the relevant asymptotic regime.
Using (20), the posterior odds after a database match are:
The DNA regime. Standard STR genotype profiling yields match probabilities on the order of $q\sim 10^{-12}$ or smaller. Even with databases of size $n\sim 10^{6}$, the product $nq\sim 10^{-6}\ll 1$ remains far below the critical scale where coincidental matches become probable.
In this extreme regime, likelihood ratios dominate the posterior odds. The prior odds $r/n$ may be small (perhaps $10^{-6}$ if we have one suspect among a million), but the likelihood ratio $s/q\sim 10^{12}$ is so enormous that posterior probabilities remain overwhelmingly high. This validates Balding’s argument: database size does not meaningfully dilute evidential weight when $nq\ll 1$.
The surveillance regime. Multi-attribute surveillance systems operate in a fundamentally different regime. With $\lambda=5$ and threshold multiplier $c=3$ (i.e., $m=15$), we have $q\sim 10^{-4}$ (Section 3.1). For populations of $n\sim 10^{6}$, the product $nq\sim 100\gg 1$ places the system far above the critical scale.
In this regime, prior odds collapse faster than likelihood ratios can compensate. Even if the likelihood ratio $s/q$ is substantial, the prior odds $r/n$ are so unfavorable that posterior probabilities remain low. Stockmarr’s caution applies: match evidence loses evidential weight as search populations grow.
Resolution. The critical scale $n_{\mathrm{crit}}\sim\exp(\lambda D)$ separates these regimes. Stockmarr and Balding are both correct in their respective contexts: DNA forensics operates where $nq\ll 1$ (likelihood-driven), while multi-attribute surveillance operates where $nq\gg 1$ (base-rate dominated). The apparent contradiction dissolves once we recognize this regime distinction.
4.1.5 Key Takeaways
  1. 1.
  2. 2.
  3. 3.
    Exponential data growth overwhelms adaptation: Reducing $q$ by a factor of $\beta$ requires increasing $k$ by only $\log(1/\beta)/(pD)$, while real-world data growth is exponential: $k(t)=k_{0}\gamma^{t}$ (Section 3.2). Thus posterior collapse is temporally inevitable.
  4. 4.
  5. 5.
Remark 4.5 (Connections to Classical Statistical Fallacies).
This section unifies the base-rate fallacy, the prosecutor’s fallacy, false discovery rate control [15], and the PPV problem in medical screening [35]. Conceptually, these phenomena are identical: all reflect Bayes’ rule under low prevalence and imperfect specificity. The widely cited argument that “most published research findings are false” [36] is the same FDR/PPV problem in another domain.
4.2 Connections to Algorithmic Fairness
The Group Dominance Effect (Theorem 3.13) has important implications for fairness in surveillance systems. When different population groups experience differential surveillance exposure, small differences in exposure rates create exponential disparities in outcomes. Proposition 3.15 shows that exposure ratios of 2–4 times generate false alert disparities exceeding 20 times near critical thresholds. This exponential amplification, which arises from Poisson tail behavior, means that even modest differences in surveillance intensity produce severe outcome inequalities.
Remark 4.6 (Structural vs. Algorithmic Bias).
The exponential amplification in part (1) is particularly striking: when both groups are screened using the same attribute set and threshold, small differences in exposure translate to exponential differences in false alert rates (Figure 1(d)).
The effect manifests temporally as different groups reach critical false-alert rates at different times: Group 2 with exposure rate $p_{2}=0.020$ fails at $k_{2}^{*}=350$ attributes, while Group 1 with $p_{1}=0.005$ remains reliable until $k_{1}^{*}=1400$. While this fourfold difference in system lifetime simply reflects the fourfold difference in exposure rates (a linear relationship), the amplification becomes exponential when comparing simultaneous false alert rates: at any intermediate $k\in(350,1400)$, Group 2 experiences exponentially more false alerts than Group 1.
Connection to algorithmic fairness. These findings relate directly to classical impossibility theorems in the algorithmic fairness literature. Kleinberg et al. [37] and Chouldechova [19] show that equalizing false positive rates, false negative rates, and calibration is impossible when base rates differ. Our analysis identifies a complementary, and more fundamental, mechanism: surveillance exposure itself creates different effective base rates ($\lambda_{g}=kp_{g}$), guaranteeing unequal false alert burdens even before any classifier is applied.
Standard fairness interventions operate at the classifier level: equalized odds [18], demographic parity, and calibration attempt to constrain predictions. None can correct the structural disparity we identify, because it arises from data collection intensity ($p_{g}$), not from how a classifier processes the collected data. Group-specific thresholds $m_{g}$ might equalize $q_{1}$ and $q_{2}$, but only by encoding exposure inequality directly. Achieving parity requires assigning a higher threshold to the more surveilled group.
This perspective connects to “fairness through awareness” [38] and formalizes “structural bias” arguments from the critical algorithm studies literature [33, 39, 22]. Disparities can be intrinsic to systems built on heterogeneous data collection, rather than artifacts of biased algorithms or training data.